solve
LN(x+3)+LNx=LN4

Answer by nerdybill(7384)

LN(x+3)+LNx=LN4
LNx(x+3)=LN4
x(x+3)=4
x^2 + 3x = 4
x^2 + 3x - 4 = 0
(x-4)(x+1) = 0
x = {-1, 4}
We can toss out the negative solution (ln of a negative does not make sense), leaving us with:
x = 4